OpenStudy (anonymous):

With interest of $1,832.00 and a principal of $16,000 for 206 days, use the ordinary interest method to determine the rate.

OpenStudy (amistre64):

A = total cost = interest + principal A = P(1+i)^n A/P = (1+i)^n n rt(A/P) = 1+i n rt(A/P) - 1= i

OpenStudy (amistre64):

if we have a daily interest rate of i, then we have a yearly rate of 360i

OpenStudy (amistre64):

well, 365i depending on what part of the world you are in